Household Consumption and Budget Constraints


Kali lives on mangoes and avocados, Pm = $5, Pa = $10, and her income is $200.


A. Identify her budget constraint equation and illustrate Kali’s budget line on a graph

with mangoes on the horizontal axis and avocados on the vertical axis.


B. Identify and illustrate on a new graph Kali’s new budget constraint if her income

doubled, the price of mangoes doubled, and the price of avocados remained constant.


C. Explain the impact on her real income when the price changed in the scenario in B

assuming income had remained constant.


1
Expert's answer
2021-12-03T08:32:29-0500

(a)

Budget constraint:

"5m+10a=200"

"m=\frac{200}{5}=40"

"a=\frac{200}{10}=20"




(b)

"10m+10a=400"

"m=\frac{400}{10}=40"

"a=\frac{400}{10}=40"

The budget constraint shifts upwards as shown:





(c)

"10m+10a=200"

"m=\frac{200}{10}=20"

"a=\frac{200}{10}=20"

If income had remained constant with the change in price, the budget constraint would have been shifted downwards as shown: